# What Should Embeddings Embed? Autoregressive Models Represent Latent Generating Distributions

### v1

@@ Abstract
Autoregressive language models have demonstrated a remarkable ability to extract latent structure from text. The embeddings from large language models have been shown to capture aspects of the syntax and semantics of language. But what should embeddings represent? We connect the autoregressive prediction objective to the idea of constructing predictive sufficient statistics to summarize the information contained in a sequence of observations, and use this connection to identify three settings where the optimal content of embeddings can be identified: independent identically distributed data, where the embedding should capture the sufficient statistics of the data; latent state models, where the embedding should encode the posterior distribution over states given the data; and discrete hypothesis spaces, where the embedding should reflect the posterior distribution over hypotheses given the data. We then conduct empirical probing studies to show that transformers encode these three kinds of latent generating distributions, and that they perform well in out-of-distribution cases and without token memorization in these settings.
@@ 1 Introduction
Autoregressive language models (LMs) are trained to predict the next token in a sequence (e.g., Bengio et al., 2000). Many large language models (LLMs) use the autoregressive objective for pretraining (e.g., Radford et al., 2019), and their document-level embeddings have been shown to capture elements of latent structure that appear in text, such as agent properties (Andreas, 2022) and syntax (Hewitt and Manning, 2019). However, it remains unclear why embeddings represent these elements, and, more formally, what information an embedding learned by a capable autoregressive predictor should capture. An understanding of what LLMs represent and why is important to the design and evaluation of architectures and optimizers, and to supporting transparency and safety.
@@
The representations formed by LLMs are challenging to analyze partly due to polysemanticity, where a neuron may activate for several distinct concepts Cunningham et al. (2023). Previous work has developed methods for probing LLM internal representations for specific concepts (Li et al., 2023; Meng et al., 2022; Zheng et al., 2023; Tenney et al., 2019). These efforts can be guided more effectively by developing a general theory of what aspects of the data embeddings should represent.
@@
In this work, we investigate several cases where the representations of autoregressive LMs can be formally connected with those of a Bayes-optimal agent. By linking the autoregressive objective to finding predictive sufficient statistics, we show that optimal content of embeddings can be identified in 1) independent identically distributed data, where the embedding should capture the sufficient statistics of the data; 2) latent state models, where the embedding should encode the posterior distribution over states given the data; and 3) discrete hypothesis spaces, where the embedding should reflect the posterior distribution over hypotheses given the data. We use probing methods to confirm that the relevant information can be decoded from LM embeddings. We also show that content that appears to be of a similar difficulty but not expected to be captured via predictive sufficient statistics is more challenging to recover from LM embeddings by probing.
Our analysis suggests that LLMs should represent latent structure that captures the posterior distribution over the generative process underlying text. Specifically, latent structure such that the next word $x_{n+1}$ is independent from previous words $x_{1:n}$ when conditioned on that structure should be recoverable from LLM embeddings. Additionally, by linking LM representations to Bayesian inference, our approach suggests that it may be effective to evaluate and interpret the behavior of LMs through comparison to more interpretable Bayes-optimal agents.
@@ 2 Related work
The embeddings produced by language models have been investigated in detail (Gupta et al., 2015; Köhn, 2015; Ettinger et al., 2016; Adi et al., 2017; Hupkes et al., 2018); for reviews, see Rogers et al. (2020) and Belinkov (2022). They have been shown to capture different aspects of the latent structure of text, including part of speech (Shi et al., 2016; Belinkov et al., 2017), syntactic number (Giulianelli et al., 2018; Conneau et al., 2018), sentence structure (Tenney et al., 2019; Hewitt and Manning, 2019; Liu et al., 2019; Lin et al., 2019), entity attributes (Gupta et al., 2015; Grand et al., 2022), sentiment (Radford et al., 2017), semantic roles (Ettinger et al., 2016; Tenney et al., 2019), world states (Li et al., 2021), and agent properties (Andreas, 2022). However, our work is motivated from a fundamentally different perspective. Instead of focusing on what is captured in the embeddings of these models, in this paper we explore why these particular kinds of structure might be represented as a consequence of statistical properties of the training data.
@@
Several previous papers have analyzed LLMs by making a connection to Bayesian inference. Of these, Xie et al. (2021), McCoy et al. (2023), and Wang et al. (2024) analyze the in-context learning behavior of LLMs. However, we study what models should encode based on the autoregressive objective that is typically used to train LLMs. Zhang et al. (2023) and Zheng et al. (2023) also connect LLM embeddings to Bayesian inference, but they focus on topic models embedded in LLMs, while we extend the connection to more general cases.
@@
Metalearned RNNs have also been shown to encode information equivalent to a Bayesian posterior distribution (Mikulik et al., 2020). Furthermore, recent work has also demonstrated that transformers behave like the Bayes-optimal predictor in linear regression settings (Panwar et al., 2024; Garg et al., 2022; Akyürek et al., 2023) and can approximate the posterior predictive distributions of probabilistic models such as Gaussian processes and Bayesian neural networks (Müller et al., 2022). We extend this analysis to general autoregressive language models and consider more general generative processes and what posterior distributions they should capture in these cases.
Figure 1: Three data generation processes where prediction of the next token $x_{n+1}$ is independent from previous tokens $x_{1:n}$ given a predictive sufficient statistic. The left corresponds to exchangeable data, the middle to latent state models, and the right to discrete hypotheses. The relevant predictive sufficient statistics are the sufficient statistic for $\theta$ (or $p(\theta|x_{1:n})$, $p(z_{n+1}|x_{1:n})$, and $p(h|x_{1:n})$ respectively). We show the embeddings learned by autoregressive transformers represent this information.
3 Three cases where optimal embeddings can be identified
@@
Assume we have a sequence $x_{1:n}$ and an autoregressive language model (LM) that predicts the next item in the sequence, $p(x_{n+1}|x_{1:n})$. We denote the LM embedding for sequence $x_{1:n}$ as $\phi_{n}=f(x_{1:n})$. The distribution $p(x_{n+1}|x_{1:n})$ is some function $g(\phi_{n})$ of this embedding, with that function implemented by the final layers of the neural network instantiating the LM. That is, the probability of the next element in the sequence $x_{n+1}$ only depends on $\phi_{n}$. This establishes our basic question: what should $\phi_{n}$ represent in order to accurately predict $x_{n+1}$?
In this section we describe three cases where this question can be answered analytically (Figure 1). First, when $x_{1:n}$ are independently sampled conditioned on an unknown parameter, $\phi_{n}$ needs only represent the sufficient statistic of this sequence. Second, when $x_{1:n}$ are generated by a state space model (in the discrete case, a hidden Markov model), $\phi_{n}$ need only represent the posterior distribution over states given $x_{1:n}$. Finally, when $x_{1:n}$ are sampled independently from one of a discrete set of latent distributions, $\phi_{n}$ need only represent the posterior distribution over hypotheses about that distribution given $x_{1:n}$. In each case we explain how $p(x_{n+1}|x_{1:n})$ factorizes to make it possible for $x_{1:n}$ to be summarized by some $\phi_{n}$ and identify the form of the corresponding $g(\phi_{n})$.
3.1 General approach
@@
The key idea behind our approach is that we can identify situations where the embedding $\phi_{n}$ contains all of the information from $x_{1:n}$ required to predict $x_{n+1}$. This idea is directly related to the notion of a sufficient statistic (Gelman et al., 2004). Given a distribution $p(x)$ with parameters $\theta$, a statistic $s(x)$ is sufficient for $\theta$ if the conditional distribution of $x$ given $s$ does not depend on $\theta$. In other words, if we only know $s$, we can estimate $\theta$ just as well as if we know the actual value of $x$. For example, if $x$ is a set of independently and identically distributed draws from a Gaussian distribution and $\theta$ is the mean of that distribution, then the mean of the sample $x$ is a sufficient statistic for $\theta$. The generative process can be rewritten to generate the sample mean $s$ given $\theta$ and then generate the variation around that sample mean independently from $\theta$, and $s$ contains all the information relevant to estimating $\theta$.
In the autoregressive setting, we care about predictive sufficiency (Bernardo and Smith, 2000). A statistic $s(x_{1:n})$ is predictive sufficient for the sequence $x_{1:n}$ if
@@
If a model performs autoregressive modeling perfectly, its embedding should represent a predictive sufficient statistic. Our three cases thus correspond to settings where predictive sufficient statistics can be easily identified and could plausibly be represented by a neural network.
@@ 3.2 Case 1: Exchangeable models
Predictive sufficiency is particularly straightforward to establish in exchangeable models, where the probability of a sequence remains the same under permutation of the order of its elements. That is, a sequence is exchangeable if $p(x_{1:N})=p(x_{\pi(1:N)})$ for some permutation $\pi$. Any exchangeable model can be re-expressed in terms of the $x_{i}$ being sampled independently and identically distributed according to a latent distribution $p(x|\theta)$ parameterized by $\theta$, with $p(x_{1:N})=\int_{\theta}\prod_{i}p(x_{i}|\theta)p(\theta)\,d\theta$ (Gelman et al., 2004). This idea leads to the following proposition:
@@
The resulting sufficient statistic also fully specifies the posterior on the parameters of the generating distribution, $p(\theta|x_{1:n})$. Sufficient statistics are easily identified for a wide range of distributions, including all exponential family distributions (Bernardo and Smith, 2000), and are easy to represent. Since the LM’s predictive distribution decomposes into the form above, this theoretical result gives us strong predictions about the contents of embeddings for models trained on exchangeable data.
3.3 Case 2: Latent state models
In a latent state model, each $x_{i}$ is generated based on a latent variable $z_{i}$. These $z_{i}$ are interdependent, with $z_{i}$ being generated from a distribution conditioned on $x_{i-1}$. Common latent state models include Kalman filters (where $x_{i}$ and $z_{i}$ are continuous and $p(x_{i}|z_{i})$ and $p(z_{i}|z_{i-1})$ are linear-Gaussian) (Kalman, 1960) and hidden Markov models (where the $z_{i}$ are discrete) (Baum and Petrie, 1966). In a latent state model, the posterior predictive distribution is
[…]
In this case, the posterior distribution $p(z_{n+1}|x_{1:n})$ captures all of the information in $x_{1:n}$ relevant to predicting $x_{n+1}$, rendering $x_{n+1}$ independent of $x_{1:n}$ when conditioned on the posterior, and acts as a predictive sufficient statistic in this model.
More formally, we can consider an embedding a representation $\phi_{n}$ such that a fixed operator $g(\phi_{n})$ can produce $p(x_{n+1}|x_{1:n})$. The posterior distribution $p(z_{n+1}|x_{1:n})$ satisfies this characterization, with Equation 2 showing that the relevant operator is the integral of $p(x_{n+1}|z_{n+1})$ over $z_{n+1}$. That operator can be easily approximated linearly and hence by a single layer of a neural network. The embedding $\phi_{n}$ thus need only represent $p(z_{n+1}|x_{1:n})$.
3.4 Case 3: Discrete hypothesis spaces
In a more specific version of Case 1, assume each $x_{i}$ is generated independently from some unknown generative model. Let $\mathcal{H}$ denote the set of hypotheses $h$ about the identity of this model. In this case, $x_{1:n}$ are exchangeable, but any sufficient statistics might be difficult to identify. The posterior predictive distribution can be written as

### v2

@@ Abstract
Autoregressive language models have demonstrated a remarkable ability to extract latent structure from text. The embeddings from large language models have been shown to capture aspects of the syntax and semantics of language. But what should embeddings represent? We show that the embeddings from autoregressive models correspond to predictive sufficient statistics. By identifying settings where the predictive sufficient statistics are interpretable distributions over latent variables, including exchangeable models and latent state models, we show that embeddings of autoregressive models encode these explainable quantities of interest. We conduct empirical probing studies to extract information from transformers about latent generating distributions. Furthermore, we show that these embeddings generalize to out-of-distribution cases, do not exhibit token memorization, and that the information we identify is more easily recovered than other related measures. Next, we extend our analysis of exchangeable models to more realistic scenarios where the predictive sufficient statistic is difficult to identify by focusing on an interpretable subcomponent of language, topics. We show that large language models encode topic mixtures inferred by latent Dirichlet allocation (LDA) in both synthetic datasets and natural corpora.
@@ 1 Introduction
Autoregressive language models (LMs) are trained to predict the next token in a sequence (e.g., Bengio et al., 2000). Many large language models (LLMs) use the autoregressive objective for pretraining (e.g., Radford et al., 2019), and their document-level embeddings have been shown to capture elements of latent structure that appear in text, such as agent properties (Andreas, 2022) and syntax (Hewitt and Manning, 2019). While it is intuitive that representing this information assists in next token prediction, we lack a formal understanding of which aspects of text should be represented and why. Given that these embeddings are often used for downstream tasks such as sequence classification and sentiment analysis, understanding what they represent and why is key to interpreting and building on these models.
@@
Analyzing the representations formed by LLMs is challenging partly due to polysemanticity, where a neuron may activate for several distinct concepts (Cunningham et al., 2023). Previous work has developed methods for probing LLM internal representations for specific concepts (Li et al., 2023; Meng et al., 2022; Zheng et al., 2023; Tenney et al., 2019). These efforts can be guided more effectively by developing a general theory of what aspects of the data embeddings should represent.
@@
In this work, we investigate several cases where the representations of autoregressive LMs can be formally connected with those of a Bayes-optimal agent. Because of their expressive architecture and their objective, LM embeddings should encode latent structure such that the next word $x_{n+1}$ is independent from previous words $x_{1:n}$ when conditioned on that structure. This corresponds to the notion of a predictive sufficient statistic. Using this connection, we show that optimal content of embeddings can be identified in two major cases: 1) independent identically distributed data, where the embedding should capture the sufficient statistics of the data; 2) latent state models, where the embedding should encode the posterior distribution over the next latent state given the data. An application of case 1 that we explore is topic models based on latent Dirichlet allocation (LDA), where the LM embedding encodes the topic mixtures of a given document. We use probing to confirm that the relevant information can be decoded from LM embeddings, and that content that is not expected to be captured via predictive sufficient statistics is more challenging to recover.
Next, we extend our analysis of exchangeable models to more realistic scenarios by generating synthetic data where only a fraction of words are generated from LDA. We confirm that language models represent topic distributions when a sufficiently large portion of the text serves this semantic component. We finally show that the encodings of pretrained LLMs (Gpt-2, Llama 2, and Bert) contain information analogous to that extracted by LDA from two natural corpora, supporting the hypothesis that LLMs implicitly perform Bayesian inference in the naturalistic setting.
Our analysis provides concrete examples where we can analytically identify what embeddings should represent, and empirical confirmation that this information is indeed captured in models trained with an autoregressive objective. By linking LM representations to Bayesian inference, our approach supports recent arguments that the behavior of LMs can be interpreted through comparison with Bayes-optimal agents (Mikulik et al., 2020). Bayesian inference encourages the agent to summarize data by constructing a generative process. Thus, understanding this connection can help LLM researchers hypothesize what features are encoded in LLM internals as done in mechanistic interpretability (Cunningham et al., 2023); by recovering latent generating distributions from LM, we show that LMs not only learn concepts, but also learn their uncertainty representation, complementing recent studies at the behavioral level (Gruver et al., 2024).
Figure 1: Three data generation processes where prediction of the next token $x_{n+1}$ is independent from previous tokens $x_{1:n}$ given a predictive sufficient statistic. The left corresponds to exchangeable data, the middle to latent state models, and the right to discrete hypotheses. The relevant predictive sufficient statistics are the sufficient statistic for $\theta$, $z_{n+1},h$ (or $p(\theta|x_{1:n})$, $p(z_{n+1}|x_{1:n})$, and $p(h|x_{1:n})$ respectively). We show the embeddings learned by autoregressive transformers represent this information.
@@ 2 Related work
The embeddings produced by language models have been investigated in detail (Gupta et al., 2015; Köhn, 2015; Ettinger et al., 2016; Adi et al., 2017; Hupkes et al., 2018); for reviews, see Rogers et al. (2020) and Belinkov (2022). They have been shown to capture aspects of the latent structure of text, including part of speech (Shi et al., 2016; Belinkov et al., 2017), sentence structure (Tenney et al., 2019; Hewitt and Manning, 2019; Liu et al., 2019; Lin et al., 2019), sentiment (Radford et al., 2017), semantic roles (Ettinger et al., 2016; Tenney et al., 2019), and agent properties (Andreas, 2022). However, our work is motivated from a fundamentally different perspective. Instead of focusing on what is captured in the embeddings of these models, we provide a general theory of language models as Bayes-optimal agents to explain why these kinds of structure might be represented due to statistical properties of the training data. Characterizing the ideal representations of a Bayes-optimal agent complements recent efforts in mechanistic interpretability (Nanda et al., 2023; Cunningham et al., 2023; Cammarata et al., 2020) by identifying optimal representations of data; this can inform the search for interpretable features that are used in LLMs’ circuit-level computations.
Implicit Bayesian inference
@@
Several previous papers have analyzed LLMs by making a connection to Bayesian inference. Of these, Xie et al. (2021), McCoy et al. (2023), and Wang et al. (2024b) analyze the in-context learning behavior of LLMs. However, we study what models should encode based on the autoregressive objective that is typically used to train LLMs. Zheng et al. (2023) focus on topic models embedded in LSTMs, while we extend the connection to more general cases. Wang et al. (2024a) provides a novel method to extract from language models concepts that come in the form of a mixed membership model (including the topic model). Our focus is more general and explanatory as we analyze why concepts that come in the form of sufficient stats should be embedded by autoregressive language models.
Belief state inference (BSI) on latent state models (Shai et al., 2024) decodes the distributions of the current latent state from the transformer embeddings on observed sequences. Our more general theory shows that this ability comes from the embedding encoding predictive sufficient statistics for the sequence, which more directly relates to the distribution of the next latent state given observations. Concurrently, Ye et al. (2024) also used de Finetti’s theorem to support probabilistic reasoning in transformers, but focused on in-context learning in the context of Bayesian linear regression. We analyze transformer internal representations and identify optimal embeddings in data that violate the exchangeability assumption.
Theory
@@
Tishby and Zaslavsky (2015) gave a theoretical analysis of optimal embeddings in deep networks from an information-theoretic perspective. We take a complementary Bayesian approach, explicitly connecting optimal embeddings to predictive sufficient statistics and presenting extensive experiments confirming that the predicted content of embeddings is tracked by transformer-based LMs. Metalearned RNNs have also been shown to encode information equivalent to a Bayesian posterior distribution (Mikulik et al., 2020). Furthermore, recent work has also demonstrated that transformers behave like the Bayes-optimal predictor in linear regression settings (Panwar et al., 2024; Garg et al., 2022; Akyürek et al., 2023) and can approximate the posterior predictive distributions of probabilistic models such as Gaussian processes and Bayesian neural networks (Müller et al., 2022). We extend this analysis to general autoregressive language models and consider more general generative processes and what posterior distributions they should capture in these cases.
3 Optimal embeddings
@@
Assume we have a sequence $x_{1:n}$ and an autoregressive language model that predicts the next item in the sequence, $p(x_{n+1}|x_{1:n})$. We denote the LM embedding for sequence $x_{1:n}$ as $\phi_{n}=f(x_{1:n})$. The distribution $p(x_{n+1}|x_{1:n})$ is a function $g(\phi_{n})$ of this embedding, with that function implemented by the final layer of the neural network instantiating the LM such that the probability of the next element $x_{n+1}$ only depends on $\phi_{n}$. This establishes our question: what should $\phi_{n}$ represent in order to accurately predict $x_{n+1}$?
@@
The key idea behind our approach is that we can identify situations where $\phi_{n}$ contains all of the information from $x_{1:n}$ required to predict $x_{n+1}$ by using the notion of a sufficient statistic (Gelman et al., 2004). Given a distribution $p(x)$ with parameters $\theta$, a statistic $s(x)$ is sufficient for $\theta$ if the conditional distribution of $x$ given $s$ does not depend on $\theta$. In other words, if we only know $s$, we can estimate $\theta$ just as well as if we know $x$. In the autoregressive setting, we care about predictive sufficiency (Bernardo and Smith, 2000). A statistic $s(x_{1:n})$ is predictive sufficient for the sequence $x_{1:n}$ if
@@
Crucially, if a model performs autoregressive modeling perfectly, it learns a predictive sufficient statistic.
In the remainder of this section we describe two cases where predictive sufficient statistics can be easily identified and could plausibly be represented by a neural network (Figure 1). These cases correspond to common applications of machine learning models as well as classic models for text such as hidden Markov models (Jurafsky and Martin, 2008). First, when $x_{1:n}$ are independently sampled conditioned on an unknown parameter, $\phi_{n}$ needs only represent the sufficient statistic of this sequence. Second, when $x_{1:n}$ are generated by a state space model (in the discrete case, a hidden Markov model), $\phi_{n}$ needs only represent the posterior distribution over the next latent state given $x_{1:n}$. In each case we explain how $p(x_{n+1}|x_{1:n})$ factorizes to make it possible for $x_{1:n}$ to be summarized by some $\phi_{n}$ and identify the form of the corresponding $g(\phi_{n})$.
@@ 3.1 Case 1: Exchangeable models
Predictive sufficiency is straightforward to establish in exchangeable models, where the probability of a sequence remains the same under permutation of the order of its elements. That is, a sequence is exchangeable if $p(x_{1:N})=p(x_{\pi(1:N)})$ for some permutation $\pi$. Any exchangeable model can be re-expressed in terms of the $x_{i}$ being sampled independently and identically distributed according to a latent distribution $p(x|\theta)$ parameterized by $\theta$, with $p(x_{1:N})=\int_{\theta}\prod_{i}p(x_{i}|\theta)p(\theta)\,d\theta$ (Gelman et al., 2004). This idea leads to the following proposition:
@@
The resulting sufficient statistic also fully specifies the posterior on the parameters of the generating distribution, $p(\theta|x_{1:n})$. Sufficient statistics are easily identified for a wide range of distributions, including all exponential family distributions (Bernardo and Smith, 2000), and are easy to represent. Since the LM’s predictive distribution decomposes into the form above, this result gives us strong predictions about the contents of embeddings for models trained on exchangeable data.
3.1.1 Case 1.1: Discrete hypothesis spaces
In a more specific version of Case 1, assume each $x_{i}$ is generated independently from some unknown generative model, and let $\mathcal{H}$ denote the discrete set of hypotheses $h$ about the identity of this model. In this case, $x_{1:n}$ are exchangeable, but any sufficient statistics might be difficult to identify. The posterior predictive distribution can be written as
@@
3.1.2 Case 1.2: Topic Models
Latent Dirichlet Allocation (LDA; Blei et al. (2001)) is an exchangeable generative model that is widely used for modelling the topic structure of documents. A document is generated from a mixture of $K$ topics. Each topic is a distribution over the vocabulary; e.g., a topic corresponding to geology might assign high probability to words such as mineral or sedimentary. We detail the full generative process in Appendix A.2.
After LDA is trained on a corpus, the inferred quantities can be used to explore the corpus. We denote by $\theta_{i}$ as the latent variable that stands for each document’s underlying topic mixture. Because topic mixture $\theta$ and word distribution $\beta$ form the sufficient statistic, our theory suggests that autoregressive LMs should implicitly encode the topic structure of a document.
3.2 Case 2: Latent state models
In a latent state model, each $x_{i}$ is generated based on a latent variable $z_{i}$. These $z_{i}$ are interdependent, with $z_{i}$ being generated from a distribution conditioned on $x_{i-1}$. Common latent state models include Kalman filters (where $x_{i}$ and $z_{i}$ are continuous and the emission and transition functions, $p(x_{i}|z_{i})$ and $p(z_{i}|z_{i-1})$, are linear-Gaussian) (Kalman, 1960) and hidden Markov models (where the $z_{i}$ are discrete) (Baum and Petrie, 1966). In a latent state model, the posterior predictive distribution is
$\hskip-5.69054ptp($ $x_{n+1}|x_{1:n})=\int p(x_{n+1}|z_{n+1})p(z_{n+1}|x_{1:n})\,dz_{n+1}.$ (2)
In this case, the posterior over the next latent state $p(z_{n+1}|x_{1:n})$ captures all of the information in $x_{1:n}$ relevant to predicting $x_{n+1}$, rendering $x_{n+1}$ independent of $x_{1:n}$ when conditioned on this statistic. If we were to drop conditional independence assumptions in a latent state model, one would need $p(x_{n+1}|z_{n+1},x_{1:n})$ instead of $p(x_{n+1}|z_{n+1})$ for the equality. Thus, $p(z_{n+1}|x_{1:n})$ is a predictive sufficient statistic in this model. However, there is another predictive sufficient statistic. We can write
$\hskip-5.69054ptp($ $x_{n+1}|x_{1:n})=\int p(x_{n+1}|z_{n})p(z_{n}|x_{1:n})\,dz_{n},$ (3)
showing $p(z_{n}|x_{1:n})$ to also be a predictive sufficient statistic. Which of these will be favored will depend on how easily the relevant integrals can be approximated.
More formally, we consider an embedding a representation $\phi_{n}$ such that a fixed operator $g(\phi_{n})$ can produce $p(x_{n+1}|x_{1:n})$. The distributions $p(z_{n+1}|x_{1:n})$ and $p(z_{n}|x_{1:n})$ satisfy this characterization, with Equations 2 and 3 showing that the relevant operator is the integral of $p(x_{n+1}|z_{n+1})$ over $z_{n+1}$ or $p(x_{n+1}|z_{n})$ over $z_{n}$ and $z_{n+1}$. That operator can be easily approximated linearly and hence by a single layer of a neural network. The embedding $\phi_{n}$ thus needs only represent $p(z_{n+1}|x_{1:n})$ or $p(z_{n}|x_{1:n})$, depending on the relative ease of approximating the relevant integrals in a specific autoregressive predictor.